OpenStudy (erinweeks):

Find a rational zero of the polynomial function and use it to find all the zeros of the function. f(x) = x^4 + 3x^3 - 5x^2 - 9x - 2

OpenStudy (erinweeks):

@chris00 can you help me with this last one

OpenStudy (anonymous):

gimmi a sec

OpenStudy (erinweeks):

okay!

OpenStudy (anonymous):

sythetic division?

OpenStudy (anonymous):

to find a rational zero we usually have to pick any random number

OpenStudy (anonymous):

so by inspection x=2 is a rational zero

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

it says find a rational zero...

OpenStudy (anonymous):

then use it to find all the other zero's

OpenStudy (anonymous):

but where'd you get 2 from?

OpenStudy (anonymous):

i said by inspection...

OpenStudy (anonymous):

they wouldn't give a nasty question...its either in the range of -3 to 3 usualy

OpenStudy (erinweeks):

@jim_thompson5910 can you help?

jimthompson5910 (jim_thompson5910):

yummydum you can use the rational root theorem to find that the possible rational roots are: -1, 1, -2, 2

OpenStudy (anonymous):

I would say to do synthetic division:

jimthompson5910 (jim_thompson5910):

ie the factors (both plus and minus) of 2, which is the last term

OpenStudy (erinweeks):

im confused..

OpenStudy (anonymous):

to do so you find the \({p\over q}'s\) which would be \(\pm1,\pm2\) \[x^3+3x^3-5x^2-9x-2\] \[\underline{2}|~~~1~~~~~~3~~~-5~~~-9~~~-2\]\[~~~~~~\underline{~~~~~~~~~2~~~~~~10~~~~~10~~~~~~~2~}\]\[~~~~~~~1~~~~~~5~~~~~~~5~~~~~~~~1~~~~~~~0\] so now you know that one of the roots are \(x=2\) and you have left \(x^3+5x^2+5x+1\) so if you divide sythetically again:\[\underline{-1}|~~~1~~~~~~~5~~~~~~~~5~~~~~~~~1\]\[~~~~~~\underline{~~~~~~~-1~~~~-4~~~~-1~}\]\[~~~~~~~~~1~~~~~~~4~~~~~~~~1~~~~~~~~0\] and now you know that another root is \(x=-1\) and you have left \(x^2+4x+1\) so that you can just plug into the quadratic equation:\[x={{b\pm\sqrt{b^2-4ac}}\over2a}\]\[x={{4\pm\sqrt{4^2-4(1)(1)}}\over2(1)}\]\[x={{4\pm\sqrt{16-4}}\over2}\]\[x={4\over2}\pm{\sqrt{12}\over2}\]\[x={2}\pm{2\sqrt{3}\over2}\]\[x=2\pm\sqrt{3}\] and so the roots of \(f(x)= x^4 + 3x^3 - 5x^2 - 9x - 2\) are: \[x={2,~~x=-1,2~~x=\sqrt{3},~~x=-2\sqrt{3}}\]

OpenStudy (anonymous):

sorry, typo: its \(x={2,~~x=-1,~~x=2\sqrt{3},~~x=-2\sqrt{3}}\)

OpenStudy (anonymous):

@ErinWeeks